A person jumps onto a cement floor from a height of 1 m and comes to rest in 0.1 second .the same person on jumping from a height of 9 m into a sand pit, comes to rest in 1 second . Compare the forces exerted on him by cement floor and sand pit.



Grade:

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Vipul

in first case

velocity v1 =√2gh1

chnage in momentum Δp1=mv1

force F1=Δp1/Δt1

in second case

velocity v2 =√2gh2

chnage in momentum Δp2=mv2

force F2=Δp2/Δt2

F1/F2=(Δt2*√h1)/(Δt1*√h2)=(1*1/0.1*3)=10/3

 

All the best.

AKASH GOYAL

AskiitiansExpert-IITD

 

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..